# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v2

@@ V Conclusion and Future Work
$\begin{split}A_{1}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Bigg\}=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using (10), (11) and (A2), we can rewrite $A_{1}$ as
@@
$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
Using (A1), the second term ($A_{2}$) of (A5) is expanded as:
$\begin{split}A_{2}&=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using the (12), (13) and (A2), we can rewrite $A_{2}$ as
@@
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A9)
Using (A3), the term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (6) can be written as:
$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A10)
Using the (A4), (A9) and (A10) in (6) gives the SINR closed-form expression in (14).
-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$.
The gradient component in (24) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
  • $0<S_{mt},I_{mt}\leq C_{1}$ (always bounded)
  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded from (25) and (26))
@@
  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{3}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{4}$,
where $C_{1}$, $C_{3}$, $C_{3}$ and $C_{4}$ are finite constants.
The partial derivatives of $g(\delta_{mi})$ is given by
$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\!\!\sum_{t=1}^{T}\!\!\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}+S_{mt}}\Bigg(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}\\ &-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\Bigg)\Bigg]-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}\\ &-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
@@
-C Derivation of closed-form SINR expression for G-PWPFZF combining
The derivation of G-PWPFZF SINR closed-form expression follows the same step as the derivation of G-PFZF SINR closed-form expression, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
The desired signal term in (6)
Using (10) and (28), then
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (C1)
Following the steps in Appendix -A, the first term in (6) can be written as
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (C2)
Similarly, the last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (C3)
Using the (C1), (C2) and (C3) in (6) gives the SINR closed-form expression in (30).

### v3

@@ V Conclusion and Future Work
$\begin{split}&A_{1}=\mathbb{E}\Big\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Big\}\\ &=\!\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}\\ &-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\\ &+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using (12) and (13), we can rewrite $A_{1}$ as
@@
$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\\ &+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A3)
The second term ($A_{2}$) of (A2) is expanded as:
$\begin{split}&A_{2}=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}\\ &-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\\ &+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
Using the (14) and (15), we can rewrite $A_{2}$ as
@@
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}\\ &+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
The term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (7) can be written as:
$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}\\ &=\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A7)
Substituting (A1), (A6) and (A7) in (7) gives the (16).
From Theorem 2, the d-LSFD weight for UE $t$ is $\mathbf{a}_{t}^{\mathrm{d}}{=}p^{u}_{t}\mathbf{D}_{t}^{-1}\mathbf{d}_{t}$. Since, $\mathbf{D}_{t}$ is a diagonal matrix, therefore, the d-LSFD for UE $t$ at AP $m$ can be written as $a_{mt}=p^{u}_{t}\frac{[\mathbf{d}_{t}]_{m}}{[\mathbf{D}_{t}]_{mm}}$. Here, the element of $\mathbf{d}_{t}$ for G-PFZF combining is given by
$\mathbf{d}_{t}]_{m}=\mathbb{E}\{\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mt}\}=\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mt}}.$
The diagonal element $\mathbf{D}_{t}]_{mm}$ is given by
$\sum\limits_{k=1}^{T}p^{u}_{k}\mathbb{E}\{|\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mk}|^{2}\}{-}p^{u}_{t}|\mathbb{E}\{\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mt}\}|^{2}+\mathbb{E}\{|\mathbf{v}_{m{i_{t}}}^{H}\mathbf{n}_{m}|^{2}\}$
$=\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!\!\!p^{{u}}_{k}\sqrt{\!(\!A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}}\!)\gamma_{mk}}{+}\!\sum\limits_{k=1}^{T}\!p^{{u}}_{k}\!(\beta_{{m}k}{-}\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k}\!){+}1,$
where $\mathbf{v}_{m{i_{t}}}=\mathbf{v}_{m{i_{t}}}^{\text{LZF}}$, if $i_{t}{\in}\mathcal{S}_{m}$, and $\mathbf{v}_{m{i_{t}}}=\mathbf{v}_{m{i_{t}}}^{\text{MR}}$, if $i_{t}{\notin}\mathcal{S}_{m}$. These expressions give the d-LSFD weights given in (18).
-E Closed-form SINR and d-LSFD expressions for G-PWPFZF combining
The derivation follows the same steps as in Appendix -D, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector. The desired signal term in (7)
Using (12) and (28), then
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (B1)
The first term in (7) can be written as
$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\\ &+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (B2)
The last term of interference in (7) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (B3)
Substituting (B1), (B2) and (B3) in (7) gives the (29).
The d-LSFD for UE $t$ at AP $m$ can be written as $a_{mt}=p^{u}_{t}\frac{[\mathbf{d}_{t}]_{m}}{[\mathbf{D}_{t}]_{mm}}$, where
$[\mathbf{d}_{t}]_{m}=\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}},$
[…]
Here, the weak pilot uses PMR combining vector insetad of MR combining vector. The above expression gives the d-LSFD weights in (30).
-F Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$. The gradient component in (26) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
  • $0<S_{mt},I_{mt},\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{1}$ (always bounded)
@@
  • $\max\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\leq L_{p}$
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{2}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{3}$,
where $C_{1}$, $C_{2}$ and $C_{3}$ are finite constants. The partial derivatives of $g(\delta_{mi})$ is given by
$\begin{split}&\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}{=}\sum_{t=1}^{T}\Bigg[\frac{\big(S_{mt}\partial_{\delta_{mi}}I_{mt}{-}I_{mt}\partial_{\delta_{mi}}S_{mt}\big)\big(I_{mt}\partial_{\delta_{mi}}S_{mt}{-}S_{mt}\partial_{\delta_{mi}}I_{mt}\big)}{(I_{mt}{+}S_{mt})^{2}I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}{+}S_{mt}}\Bigg(\frac{I_{mt}\partial^{2}_{\delta_{mi}}S_{mt}-S_{mt}\partial^{2}_{\delta_{mi}}I_{mt}}{I_{mt}^{2}}\\ &-\frac{2\,\partial_{\delta_{mi}}I_{mt}\big(I_{mt}\partial_{\delta_{mi}}S_{mt}{-}S_{mt}\partial_{\delta_{mi}}I_{mt}\big)}{I_{mt}^{3}}\Bigg)\Bigg]{-}\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}{-}\delta_{mi}^{2}>0\}}\big((1{-}2\delta_{mi})^{2}{-}2(\delta_{mi}{-}\delta_{mi}^{2})\big){+}2\lambda_{2}\,\mathbf{1}_{\{2{-}A{>}0\}}\Bigg].\end{split}$ (C1)